OpenStudy (anonymous):

Find the coordinates of the points that are 10 units away from the origin and have a y-coordinate equal to -6. (smallest x value, and largest x value)

OpenStudy (mertsj):

\[\sqrt{(x-0)^2+(-6-0)^2}=10\] \[x^2+36=100\] \[x^2=64\] \[x=\pm8\]

OpenStudy (anonymous):

thank you!!

OpenStudy (mertsj):

yw